OpenStudy (anonymous):

Create a quadratic function, f(x), in vertex form. The a should be between 4 and –4, the h will be your birth month, and the k will be your birth day. Write your equation below.

OpenStudy (anonymous):

What do you want a, h, and k to be?

OpenStudy (anonymous):

H=1 K=29 A=2

OpenStudy (anonymous):

f(x)=a(x-h)^2 + k

OpenStudy (anonymous):

So, f(x)=2(x-1)^2 + 29

OpenStudy (anonymous):

Do you know how to FOIL?

OpenStudy (anonymous):

ohhh okay got it! and by any chance do you know how to convert it into standard form? can you like explain to me how

OpenStudy (anonymous):

First you have to simplify it so FOIL it out and combine any like terms.

OpenStudy (anonymous):

f(x)=2(x-1)(x-1)+29 f(x)=2(x^2-2x+1)+29 f(x)=2x^2-4x+2+29 f(x)=2x^2-4x+31

OpenStudy (anonymous):

thank you for your help

OpenStudy (anonymous):

So in standard form it would be 0=2x^2-4x+31 OR 2x^2-4x=31